40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 13 02:52:06 2008, Lisa, Configuration, LSC, Locking with 3f 
    Reply  Thu Jul 24 01:04:01 2008, rob, Configuration, LSC, IFR2023A (aka MARCONI) settings 
Message ID: 373     Entry time: Thu Mar 13 02:52:06 2008     Reply to this: 729
Author: Lisa 
Type: Configuration 
Category: LSC 
Subject: Locking with 3f 
Today we have tried to use the reflected signal demodulated at 3*f1 ~ 99 MHz (REFL31) for length control.
This signal is cool because it is generated by the beating of sidebands, so it is not very sensitive to what the carrier does inside the IFO.
In particular, its gain and the demodulation phase shouldn't change much while changing the CARM offset during the locking sequence.
The idea is therefore to use REFL31_I and REFL31_Q for controlling MICH and PRCL, with the goal of making the lock acquisition sequence more robust.

We minimized hardware changes by using the 199MHz demodulation board, changing the local oscillator to 99.586317 MHz, with an amplitude of +10 dbm (the 3f signals are therefore acquired as LSC-PD6_I and LSC-PD6-Q).

We locked both the PRM and the DRM in a stable way using the REFL31_I and REFL31_Q, after tuning the demodulation phase (50) and removing their offsets.
On the other hand, we weren't able to acquire the lock in the DRM configuration directly by using the 3f signals. We needed instead to use the f signals first, and switch to the 3f signals once the lock was already acquired, otherwise ending up locking DRM at a different working point.
One explanation for that might be the fact that the beam impinging upon the 3f diode is too big compared with the diode size (only 1 mm, half of the size of the f1 diode).
For these reason, in presence of misalignments, some of the reflected light goes in high order modes, which can be partially (or all) off the diode, thereby generating multi-zero crossing in the demodulated error signal.

The next step before making the test with the whole IFO is therefore to modify the telescope in front of the 3f diode in order to reduce the beam size and repeat the tests we did tonight in DRM configuration.


P.S.: We made a test by changing the frequency of the local oscillator by a little bit and then coming back to the original value. We observed that the phase of the signal can change, so every time this frequency is moved the 3f demod phase need to be retuned.

John, Rob, Rana, Lisa
ELOG V3.1.3-